If the density of water is 1,00 g/cm3 and atmospheric pressure is 1,013 × 105 Pa, the depth in water (in m) where the absolute pressure is four times atmospheric pressure, is equal to


1
Expert's answer
2022-06-06T08:54:32-0400

"\frac{\rho gh}p+1=4,"

"h=\frac{3p}{\rho g}=31~m."